The sum of the ages of afraz and Sam is 33 years. In two years, 3 times Afraz's age will be one year less than 4 times sams age. How old is each now?
Answer by Boreal(15235)

A+S=33
3(A+2)+1=4(S+2). which means that if I add 1 to 3 times A's age in two years, I get 4 times S. age in 2 years.
3A+7=4S+8
But A=33-S
substitute
99-3S+7=4S+8
98=7S
S=14 now
A=19 now
In two years A will be 21 and 3 times that is 63. That is 1 year less than 4 times S. age, which is 64.